Circular No. 8658

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 YQ_{127} (LINEAR) A number of reports have been received describing the cometary nature

of 2005 YQ_{127}, an object discovered by LINEAR and presumed asteroidal. The discovery position (originally published on MPS 156789) is given below, while additional astrometry and a revised orbit (T = 2005 Nov. 3.96 TT, q = 1.9202 AU, e = 0.5042, P = 7.62 yr, ω = 152^{o}.32, Ω = 276^{o}.37, i = 16^{o}.74, equinox 2000.0) were published on MPEC 2006-A57. M. Mannucci, N. Montigiani, and S. Riccetti (Florence, Italy, 0.25-m reflector + CCD) reported that images taken on Jan. 6.89 and 9.89 UT show a diffuse coma.

2005 UT  α_{2000}  δ_{2000}  Mag.
Dec. 28.28042  5^{h}58^{m}03.^{s}15  +25^{o}25′49′′.5  18.0

SUPERNOVAE 2006D, 2006E, 2006F Three apparent supernovae have been reported from unfiltered CCD

images: 2006D by C. Colesanti, C. Jacques, E. Pimentel, and T. Napoleao (BRASS survey; cf. IAUC 8589); 2006E independently by three survey programs [by T. Puckett and V. Reddy (0.60-m automated supernova patrol telescope; cf. IAUC 8657; source of tabulated data below); by M. Baek and W. Li (LOSS/KAIT; cf. IAUC 8657); and by M. Migliardi (Cortina, Italy, 0.5-m “Ullrich” telescope in the course of the CROSS program; cf. IAUC 7373; communicated via A. Dimai, Cortina d’Ampezzo, Italy)]; and 2006F by M. Villi (Piacenza, Italy, also with the CROSS program, and via Dimai).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006D Jan. 11.19 12^{h}52^{m}33.^{s}94 − 9^{o}46′30′′.8 15.6 13′′ W, 5′′.5 N
2006E Jan. 12.45 13 53 28.65 + 5 12 22.8 14.3 31′′.1 E, 6′′.3 S
2006F Jan. 11.78 2 28 11.37 +19 36 13.5 17.3 3′′ E, 15′′ N

Additional unfiltered CCD magnitudes of 2006D in MCG –01-33-34: 2005 July 7.85 UT, [19.0; 2006 Jan. 13.12, 15.5. Also, nothing is visible at the location of 2006D on a red Palomar Survey plate from 1956. Additional magnitudes of 2006E in NGC 5338: 2004 Jan. 23, [19.8 (Puckett); 2005 June 26.25, [19.0 (KAIT); 2006 Jan. 13.15, 14.2 (Migliardi); 13.52, 14.3 (Puckett, remotely with a 0.35-m reflector at Portal, AZ); 13.58, 14.3 (KAIT). Li provides position end figures 28^{s}.52, 22′′.2 for 2006E; Migliardi provides position end figures 28^{s}.52, 22′′.3. Additional approximate CCD magnitudes of 2006F in NGC 935: 2005 Nov. 21, [18.5; 2006 Jan. 13.76, 16.7. SN 2006F is also absent on Palomar Sky Survey infrared, red, and blue plates.

2006 January 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green